OpenStudy (anonymous):

In a 30-60-90 triangle, the shorter leg has length of 8√3 m. Find the length of the other leg (L) and the hypotenuse (H).

OpenStudy (anonymous):

ratio in 30 -60-90 triangle are \(1:\sqrt{3}:2\)

OpenStudy (anonymous):

so short side is \(8\sqrt{3}\) then the long side is \(8\sqrt{3}\times \sqrt{3}=8\times 3=24\) and the hypotenuse is double the short side